OpenStudy (anonymous):

Could someone please help me with this? I need all the steps too, having a hard time understading... 4x+5y=13 3x+y=-4 and 1/3x+1/2y=5 x-2/3y=1 For the 1st one, to I multiply 2nd row by -5 to eliminate y? And in the 2nd, to I muliply everything by 3 in 2nd row to get rid of denominator?

OpenStudy (chaise):

4x+5y=13 3x+y=-4 To solve these using elimination multiply every term in the bottom equation by 5. 4x+5y=13 15x+5y=-20 We can now eliminate y by subtracting the first equation by the second. (although it doesn't matter which way you do it. 4x+5y=13 -9x = 33 x = 33/-9 x = -3 and 2/3 We can substitute this back into the original equation to solve for x. 4x+5y=13 4(-3 and 2/3) + 5y = 13 -44/3 + 5y = 13 5y = 88/3 y = 88/3/5 y = 83/15 y = 5.5333 x = -3 and 2/3 y = 5.5333 or 83/15 I'm 90% sure this is right, sorry if it isn't (: